A circle passes through the points (-1,3)and (5,11) and its radius is 5. Then its centre is?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear tadi parthi

Let the center is (h,k)

so  (h+1)2 + (k-3)2 = (h-5)2 +(k-11)2 =52

 so (h+1)2 + (k-3)2 = (h-5)2 +(k-11)2

simplify

3h +4k=34 ...............1

 

and (h+1)2 + (k-3)2 =52

      h2 +k2 +2h -6k  =15 ..............2

solve 1st and 2nd equation

that will desired solution

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin